After a late-winter freeze in Florida, the supply curve for early-season blueberries was

estimated as: QS = - 500 + 5 000 P.The demand curve for blueberries is:

QD = 19 000 – 1 500 P

P: measured in dollars, Q: measured in pints

1) Before the freeze, the equilibrium price was $0.50 per pint. Find the equilibrium P and Q after the freeze

2) There is a price ceiling in place of $1.00 per pint. If there is no illegal trading, how many pints of blueberries will be sold?



1
Expert's answer
2021-12-03T11:54:50-0500

(1)

"Q_S=-500+5000P"

"Q_D=19000-1500P"

At equilibrium:

"Q_S=Q_D"

"-500+5000P=19000-1500P"

"P=\frac{19500}{6500}=3"

"Q=-500+5000(3)=14500"


(2)

"Q=-500+5000(1)"

"Q=4500"